# with Knowledge Distillation for Question Answering

### v4

@@ Appendix I M2QA-Bench Examples
Regarding the effect of modalities, results from Tables 9 and 10 suggest that the modalities themselves are not the most critical factor. Instead, the complexity of how the question integrates the modalities plays a more significant role. For M2QA-Bench, models performed better when visual understanding was not required, with tables and text being the primary contributors to the results. In contrast, for MultiModalQA, models tended to perform better on image-based questions, highlighting the importance of how the question leverages the modalities. For questions involving both modalities, smaller models struggled more to produce correct answers, while larger models performed better in terms of EM. It is important to note, however, that due to the substantial difference in the number of samples containing both image and table modalities compared to those with only one modality, the reported results are not directly comparable. Refer to Figures 13 and 14 for the outputs of different model families fine-tuned on either real or synthesized data. Moreover, Figure 15 shows outputs from different model sizes within the same family, fine-tuned on either real or synthesized data.
Appendix K Synthesizing Data vs. Paraphrasing Exisiting Human Annotated Datasets
Figure 12: hey
Paraphrasing questions from existing datasets introduces surface-level linguistic changes but preserves the original semantic intent and reasoning pathways, offering only marginal improvements in model training. In contrast, the data synthesized by FM2DSis intentionally crafted to introduce diverse question structures, span multiple domains, and require varied types of reasoning, pushing models toward more comprehensive multimodal understanding. To compare these approaches, we trained InternVL-2-8B using 1k samples from three settings: (I) the original MultimodalQA dataset, (II) a paraphrased version of MultimodalQA where questions were reworded using GPT-4o with the prompt ”Please paraphrase the following question: [Question]” and (III) synthesized samples generated by FM2DS. For all conditions, we used full Wikipedia documents as sources. Figure 12 presents the results, showing that while paraphrasing provides a slight improvement, synthesizing new, high-quality samples with FM2DS leads to a substantial performance gain.
Appendix L Statistics on Usages of Each Validation Stage of FM2DS
As described in Section 3 and illustrated in Figure 2, FM2DS incorporates multiple validation stages to enhance data quality. It is essential to analyze how frequently each stage rejects the initially generated outputs. Table 12 presents statistics based on generating 1,000 examples using GPT-4o. Among the stages, question validation has the highest rejection rate, suggesting that this step is the most challenging. This may be because generating a question requires the model to synthesize all relevant knowledge and fully grasp the context. In contrast, answer validation benefits from the guidance provided by the question, making the task relatively easier. Query validation appears to be even more straightforward, as it primarily involves formatting the reasoning steps—something the model has effectively done during answer generation. Additionally, the use of question-specific image captions during answer generation likely contributes to a lower error rate by helping the model locate the correct information only text modality.
Stage Average Rejection Rate Average Rejection Rate
Question Validation 131 11.58%
Answer Validation 76 7.06%
Query Validation 58 5.48%
@@
Figure 13: Analysis of model responses to the question: ”Which artist released an album in December 1969 featuring a record on its cover?” from MultimodalQA dataset reveals that fine-tuning on FM2DS eliminates hallucination (marked by the confused robot sign) seen in model fine-tuned on real data. This example highlights how fine-tuning improves reasoning by aligning the model’s answers with both visual and textual evidence.
@@
Figure 14: Analysis of model responses to the question: ”Which street was paved with boards; Little Champlain Street, Quebec City, 1916 or Quebec City Rue Saint-Louis winter 2010?” from the WebQA dataset demonstrates that fine-tuning with FM2DS data effectively eliminates hallucination (indicated by the confused robot sign). This example underscores fine-tuning with FM2DS-generated data improves the model’s focus on fine-grained visual details relevant to the question. Here, InternVL-2-8B fine-tuned on real data hallucinated but reached the correct answer using its pre-trained knowledge.
@@
Figure 15: Responses from InternVL-2 models of various sizes (8B, 26B, 40B, and 76B) to the question: ”In the most important project of Qin Shi Huang, what geometric shape was used in the watchtowers when viewed from inside?” from M2QA-Bench illustrate that in examples like this, which requires detailed visual understanding, smaller models often hallucinate, providing inconsistent answers (e.g., square, rectangle) without grounding in the provided document. Larger models, however, perform better on this task and have less hallucination.

### v5

@@ Appendix I Additional Statistics & Information on M2QA-Bench
Regarding the effect of modalities, results from Tables 13 and 14 suggest that the modalities themselves are not the most critical factor. Instead, the complexity of how the question integrates the modalities plays a more significant role. For M2QA-Bench, models performed better when visual understanding was not required, with tables and text being the primary contributors to the results. In contrast, for MultiModalQA, models tended to perform better on image-based questions, highlighting the importance of how the question leverages the modalities. For questions involving both modalities, smaller models struggled more to produce correct answers, while larger models performed better in terms of EM. It is important to note, however, that due to the substantial difference in the number of samples containing both image and table modalities compared to those with only one modality, the reported results are not directly comparable. Refer to Figures 13 and 14 for the outputs of different model families fine-tuned on either real or synthesized data. Moreover, Figure 15 shows outputs from different model sizes within the same family, fine-tuned on either real or synthesized data.
@@
Figure 13: Analysis of model responses to the question: "Which artist released an album in December 1969 featuring a record on its cover?" from MultimodalQA dataset reveals that fine-tuning on FM2DS eliminates hallucination (marked by the confused robot sign) seen in model fine-tuned on real data. This example highlights how fine-tuning improves reasoning by aligning the model’s answers with both visual and textual evidence.
@@
Figure 14: Analysis of model responses to the question: "Which street was paved with boards; Little Champlain Street, Quebec City, 1916 or Quebec City Rue Saint-Louis winter 2010?" from the WebQA dataset demonstrates that fine-tuning with FM2DS data effectively eliminates hallucination (indicated by the confused robot sign). This example underscores fine-tuning with FM2DS-generated data improves the model’s focus on fine-grained visual details relevant to the question. Here, InternVL-2-8B fine-tuned on real data hallucinated but reached the correct answer using its pre-trained knowledge.
@@
Figure 15: Responses from InternVL-2 models of various sizes (8B, 26B, 40B, and 76B) to the question: "In the most important project of Qin Shi Huang, what geometric shape was used in the watchtowers when viewed from inside?" from M2QA-Bench illustrate that in examples like this, which requires detailed visual understanding, smaller models often hallucinate, providing inconsistent answers (e.g., square, rectangle) without grounding in the provided document. Larger models, however, perform better on this task and have less hallucination.